Please show me the process to use to work this problem using the distribution property process. thanks for your help!
Michael McGinnis invested some money at 7% and $4000 less than this amount at 5%. He earned $520 interest the first year. How much did he invest at 5%?

Answer by checkley75(3666)

.07x+.05(x-4000)=520 [7% of x plus 5% of (4000-x)]=520 interest
.07x+.05x-200=520
.12x=520+200
.12x=720
x=720/.12
x=6000 invested @ 7%
6000-4000=2000 invested @ 5%
proof
.07*6000+.05*2000=520
420+100=520
520=520